FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE B′ ^{2}Σ^{+} - X ^{2}Σ^{+} TRANSITIONS OF MgH AND MgD

\underline{A. SHAYESTEH}, K. TERESZCHUK, and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, ON, N2L 3G1, Canada..

Emission spectra of MgH and MgD have been investigated in the 8000 - 22000 cm^{−1} region using a high resolution Fourier transform spectrometer. These molecules were generated in a furnace-discharge source, at about 900 K and 333 mA discharge current with magnesium and a mixture of Ar and H_{2} or D_{2} gases. The recorded spectra contain not only the well-known A ^{2}Π - X ^{2}Σ^{+} transitions, but also the B′ ^{2}Σ^{+} - X ^{2}Σ^{+} transitions of MgH and MgD. We obtained data for v′′ = 2 to 9 for the ground state of MgH and v′′ = 3 to 13 for MgD. Analyses of the data is in progress, and will lead to improved potential energy curves for the X ^{2}Σ^{+} states of MgH and MgD. These results will be presented at the time of symposium.